# BeautyGuard: Designing a Multi-Agent Roundtable System for Proactive Beauty Tech Compliance through Stakeholder Collaboration

### v1

@@ 1. Introduction
Despite routine review rituals, current practice shows three recurring pain points. First, fragmentation and inconsistency: checklists live in different teams, versions drift, and outcomes depend on who is in the room. Second, thin precedent access: reviewers struggle to assemble internal cases and external actions into a traceable evidence chain. Third, weak support for mitigation: many processes stop at compliance verdicts, while beauty scenarios often require concrete ways to lower risk (e.g., consent redesign, field minimization, anonymization/watermarking, audience controls). Foundational HCI research has documented similar gaps, Holstein et al. (2019) showed that fairness and accountability practices in industry are often ad-hoc and poorly aligned with available tools, while Madaio et al. (2020) demonstrated that checklist-based practices benefit from co-design with practitioners to ensure workflow fit. Governance itself is sociotechnical, requiring alignment with organizational processes and roles rather than model-only fixes (Feng et al., 2025). At the same time, LLM reliability issues such as hallucination demand auditable reasoning and source grounding in compliance settings (Huang et al., 2025; Sovrano et al., 2025).
We propose a stakeholder-informed multi-agent LLM roundtable that mirrors enterprise roles: a Legal Interpreter for policy reading, a Rule Checker for item-by-item scanning, a Precedent Researcher for internal/external case retrieval, and a Risk Planner for graded mitigation. Multi-agent research shows that role decomposition, critique, and consensus can improve reasoning and planning over single models (Park et al., 2023; Du et al., 2024). Our system has each agent produce traceable, role-specific judgments that are then synthesized into a structured plan, shifting from rejection-based review to proactive risk paths aligned with how cross-functional teams actually work.
We studied this approach with six experts (two legal, four IT) in a multinational beauty enterprise, identified challenges in current reviews, built a working prototype, and evaluated it with SUS (Bangor et al., 2008), NASA-TLX (Hart and Staveland, 1988), and interviews. The goal is not to replace experts, but to standardize early analysis, surface alternatives, and link recommendations to rules and precedents.
@@ 2.1. Compliance Challenges and Emerging Practices in the Beauty Tech Industry
Beauty tech applications increasingly rely on LLMs for customer interaction and content generation, but these systems surface unique compliance risks. Studies show that AI-generated product descriptions in cosmetics and fashion often embed gender stereotypes, using different persuasive language and assumptions for “men’s” versus “women’s” items (Kelly et al., 2025). Such systematic bias reflects broader concerns about fairness and inclusivity in algorithmically mediated beauty platforms.
Beyond bias, generative AI introduces regulatory challenges for advertising, data use, and cultural sensitivity. Outputs must align with diverse standards across regions, from advertising law to hate speech regulation (Hacker et al., 2023). Because many beauty systems process sensitive biometric data, compliance often requires explicit consent, opt-out options for profiling, and safeguards for cross-border data transfers. Public response to face-based “beauty detection” technologies underscores these concerns, with studies documenting widespread unease and accusations of pseudoscience (Ullstein et al., 2024).
In response, emerging practices seek to embed compliance into system design. Privacy-preserving machine learning can reduce exposure to raw face data, while adversarial computer vision methods enable protective filters against unwanted recognition (Sun et al., 2024). Legal analyses further highlight ambiguity in how the EU AI Act classifies beauty-related systems as “high-risk,” complicating the adoption of harmonized standards (Golpayegani et al., 2023). These uncertainties suggest that beyond technical fixes, organizations need structured methods to surface risks early, reconcile multiple regulatory perspectives, and plan compliance strategies, an open challenge our multi-agent roundtable system directly addresses.
@@ 2.2. Multi-Agent Systems for Enterprise Decision Support
Multi-agent LLM architectures organize specialized models into role-based teams to solve complex tasks more reliably than single, monolithic models. Recent engineering and empirical work demonstrates clear benefits of role decomposition (e.g., planner/coder, reviewer/implementer) across domains such as collaborative code generation and multi-robot control (Hong et al., 2024; Qian and Cong, 2023; Mandi et al., 2024; Zhang et al., 2023). Tooling efforts make these designs accessible: AutoGen provides primitives for asynchronous agent messaging and task decomposition, enabling developers to compose agent societies for practical workflows (Wu et al., 2024).
Multi-agent designs also improve reasoning through iterative critique and consensus mechanisms. Studies show that agent ensembles that debate, critique, and vote yield higher factuality and problem-solving accuracy than single models, especially on multi-step reasoning tasks (Du et al., 2024, 2023). Generative agents augmented with memory and reflection further extend this capability by maintaining long-term state and producing contextually coherent plans over time (Park et al., 2023). Empirical work indicates that such multi-agent interactions can produce emergent coordination and rudimentary theory-of-mind behavior, although limitations remain for long-horizon planning and hallucination control (Li et al., 2023; Xiao et al., 2023).
Despite rapid progress, prior work primarily targets open-ended problem solving (coding, games, simulation) or standalone agent benchmarks; few studies situate multi-agent LLMs within enterprise decision workflows that require legal accountability, cross-functional interpretation, and auditable reasoning (Zhang et al., 2024; Guo et al., 2024). Existing agent teams typically seek a single correct output, whereas organizational decisions often demand multiple perspectives, traceable rationales, and paths for mitigation when compliance risks are identified.
@@ 2.3. Human-AI Collaboration for Compliance Review
Regulatory compliance is a high-stakes, knowledge-intensive process traditionally handled by legal and domain experts. Recent work emphasizes that AI governance is sociotechnical: effective systems must integrate organizational workflows, stakeholder perspectives, and human expertise rather than focusing solely on technical accuracy (Feng et al., 2025). Designing such systems also extends beyond interface design to model behavior and training data, requiring close collaboration between UX, engineering, and domain experts (Subramonyom et al., 2022).
AI tools have begun to assist in compliance-related work, but with mixed results. Studies show that GPT-4 can help developers prepare documentation for the EU AI Act (Sovrano et al., 2025) by identifying gaps, though outputs were often incomplete or irrelevant and required expert intervention . Similarly, LLM-based assistants in legal practice can support case research or document review, yet even with retrieval augmentation, hallucination remains nontrivial (17–33%) (Magesh et al., 2025). These results suggest AI can reduce workload, but expert oversight is indispensable.
@@ 2.4. Generative Agents for Critical Thinking and Advice
Generative agents extend LLM capabilities by storing experiences, reflecting on past interactions, and using these reflections to guide future planning. Such mechanisms improve task performance and produce more coherent long-term behavior (Park et al., 2023; Shinn et al., 2023). Beyond synthetic tasks, reflective techniques have also been adapted for coaching scenarios: a GPT-4-based chatbot using motivational interviewing improved users’ readiness for behavior change while mitigating potential harms through dialogue design (Meyer and Elsweiler, 2025). These findings highlight the potential of agents to act as cognitive partners that encourage reflection rather than dispensing answers.
At the same time, reflective support raises new cognitive and interaction challenges. Studies emphasize that users collaborating with generative tools must develop self-awareness of their goals, calibrate confidence in AI capabilities, and flexibly adapt their strategies (Tankelevitch et al., 2024). Experimental evidence further shows that the design of AI assistance profoundly shapes cognition and motivation: when AI generated ideas all at once, participants invested more effort in synthesizing suggestions, whereas incremental support led to higher confidence in their own ideas (Xu et al., 2025). These results suggest that how AI support is structured directly influences user agency.
Building on this literature, recent work calls for positioning AI contributions to complement rather than dominate human reasoning (Reicherts et al., 2025). Yet prior studies remain focused on general creativity and coaching contexts, without addressing enterprise compliance workflows where reflective synthesis must reconcile legal, technical, and managerial perspectives. Our system fills this gap by embedding reflective generative agents into a roundtable framework, encouraging critical thinking and multi-perspective planning in regulated industry settings.
@@ 7.4. Data Collection
Quantitative Measures. We collected standardized usability and workload metrics through validated instruments. SUS (Bangor et al., 2008) provided 10-item Likert scale measurements (1-5) of system usability, with all questions rephrased as positive statements to ensure consistent interpretation. NASA-TLX (Hart and Staveland, 1988) assessed cognitive workload across three key dimensions (Mental Demand, Effort, and Frustration) using the same 5-point scale format where higher scores indicate lower workload. Additionally, we measured trust and adoption metrics through custom 5-point Likert scales covering: Trust in Output, Usage Intention, Workflow Integration, and Explanation Clarity.
@@ 7.5. Data Analysis
Qualitative interview data were analyzed using thematic analysis (Braun and Clarke, 2006), with two researchers independently coding transcripts to identify recurring themes related to system effectiveness, workflow integration, and areas for improvement.
@@ 9. Discussion
This study tackled the specific challenge of enterprise compliance review in the beauty tech industry, where AI-driven products face unique regulatory complexities around biometric data handling, skin-tone fairness, and cross-border compliance (Kelly et al., 2025; Ullstein et al., 2024). We addressed three research questions: 1) current practices and challenges in beauty tech compliance review, 2) the design of a multi-agent LLM system to address these challenges, and 3) the effectiveness and user perception of such a system. We conducted a formative study with six experts (two legal, four IT) to identify domain-specific challenges. Based on these findings, we developed BeautyGuard, a multi-agent roundtable system that provides automated pre-review, role-specialized analysis, and graded risk mitigation. Six participants evaluated BeautyGuard using standardized metrics (SUS, NASA-TLX). The results showed that BeautyGuard achieved exceptional usability (SUS: 77.5/100), minimal workload across all dimensions.
While our focus was beauty tech compliance, the underlying challenges—fragmented pre-review processes, departmental silos, and binary decision-making—are common across regulated industries (Feng et al., 2025; Subramonyom et al., 2022). This suggests that our multi-agent roundtable approach has broader applicability beyond the beauty domain, with potential applications in healthcare, financial services, and other regulated sectors that require coordinated expert review processes.
@@ 9.3. How to generalize the current approach from beauty tech to other regulated industries?
Our beauty tech focus revealed domain-specific challenges—biometric data handling, skin-tone fairness, and cultural sensitivity—that required specialized compliance knowledge (Kelly et al., 2025; Ullstein et al., 2024). However, the underlying organizational challenges we identified are prevalent across regulated industries. Multi-agent systems have shown effectiveness in coordinating diverse expertise for complex decision-making tasks (Hong et al., 2024; Qian and Cong, 2023), and our results demonstrate this applies to enterprise compliance contexts.
Similar principles can be generalized to other regulated domains since core challenges—fragmented processes, departmental silos, and binary decision-making—are documented across healthcare (Subramonyom et al., 2022), financial services, and other high-stakes industries (Feng et al., 2025). The fundamental multi-agent architecture and roundtable coordination model remain applicable, with only domain-specific knowledge bases and regulatory mappings requiring adaptation.

### v2

@@ 1. Introduction
Despite routine review rituals, current practice shows three recurring pain points. First, fragmentation and inconsistency: checklists live in different teams, versions drift, and outcomes depend on who is in the room. Second, thin precedent access: reviewers struggle to assemble internal cases and external actions into a traceable evidence chain. Third, weak support for mitigation: many processes stop at compliance verdicts, while beauty scenarios often require concrete ways to lower risk (e.g., consent redesign, field minimization, anonymization/watermarking, audience controls). Foundational HCI research has documented similar gaps, 10.1145/3290605.3300830 showed that fairness and accountability practices in industry are often ad-hoc and poorly aligned with available tools, while 10.1145/3313831.3376445 demonstrated that checklist-based practices benefit from co-design with practitioners to ensure workflow fit. Governance itself is sociotechnical, requiring alignment with organizational processes and roles rather than model-only fixes (10.1145/3706599.3706747). At the same time, LLM reliability issues such as hallucination demand auditable reasoning and source grounding in compliance settings (10.1145/3703155; sovrano2025simplifying).
We propose a stakeholder-informed multi-agent LLM roundtable that mirrors enterprise roles: a Legal Interpreter for policy reading, a Rule Checker for item-by-item scanning, a Precedent Researcher for internal/external case retrieval, and a Risk Planner for graded mitigation. Multi-agent research shows that role decomposition, critique, and consensus can improve reasoning and planning over single models (park2023generative; pmlr-v235-du24e). Our system has each agent produce traceable, role-specific judgments that are then synthesized into a structured plan, shifting from rejection-based review to proactive risk paths aligned with how cross-functional teams actually work.
We studied this approach with six experts (two legal, four IT) in a multinational beauty enterprise, identified challenges in current reviews, built a working prototype, and evaluated it with SUS (bangor2008empirical), NASA-TLX (hart1988development), and interviews. The goal is not to replace experts, but to standardize early analysis, surface alternatives, and link recommendations to rules and precedents.
@@ 2.1. Compliance Challenges and Emerging Practices in the Beauty Tech Industry
Beauty tech applications increasingly rely on LLMs for customer interaction and content generation, but these systems surface unique compliance risks. Studies show that AI-generated product descriptions in cosmetics and fashion often embed gender stereotypes, using different persuasive language and assumptions for “men’s” versus “women’s” items (Kelly_2025). Such systematic bias reflects broader concerns about fairness and inclusivity in algorithmically mediated beauty platforms.
Beyond bias, generative AI introduces regulatory challenges for advertising, data use, and cultural sensitivity. Outputs must align with diverse standards across regions, from advertising law to hate speech regulation (10.1145/3593013.3594067). Because many beauty systems process sensitive biometric data, compliance often requires explicit consent, opt-out options for profiling, and safeguards for cross-border data transfers. Public response to face-based “beauty detection” technologies underscores these concerns, with studies documenting widespread unease and accusations of pseudoscience (10.1145/3630106.3659038).
In response, emerging practices seek to embed compliance into system design. Privacy-preserving machine learning can reduce exposure to raw face data, while adversarial computer vision methods enable protective filters against unwanted recognition (Sun_2024_CVPR). Legal analyses further highlight ambiguity in how the EU AI Act classifies beauty-related systems as “high-risk,” complicating the adoption of harmonized standards (10.1145/3593013.3594050). These uncertainties suggest that beyond technical fixes, organizations need structured methods to surface risks early, reconcile multiple regulatory perspectives, and plan compliance strategies, an open challenge our multi-agent roundtable system directly addresses.
@@ 2.2. Multi-Agent Systems for Enterprise Decision Support
Multi-agent LLM architectures organize specialized models into role-based teams to solve complex tasks more reliably than single, monolithic models. Recent engineering and empirical work demonstrates clear benefits of role decomposition (e.g., planner/coder, reviewer/implementer) across domains such as collaborative code generation and multi-robot control (hong2024metagpt; qian2023communicative; mandi2024roco; zhang2023building). Tooling efforts make these designs accessible: AutoGen provides primitives for asynchronous agent messaging and task decomposition, enabling developers to compose agent societies for practical workflows (wu2024autogen).
Multi-agent designs also improve reasoning through iterative critique and consensus mechanisms. Studies show that agent ensembles that debate, critique, and vote yield higher factuality and problem-solving accuracy than single models, especially on multi-step reasoning tasks (pmlr-v235-du24e; du2023improving). Generative agents augmented with memory and reflection further extend this capability by maintaining long-term state and producing contextually coherent plans over time (park2023generative). Empirical work indicates that such multi-agent interactions can produce emergent coordination and rudimentary theory-of-mind behavior, although limitations remain for long-horizon planning and hallucination control (li-etal-2023-theory; xiao2023simulating).
Despite rapid progress, prior work primarily targets open-ended problem solving (coding, games, simulation) or standalone agent benchmarks; few studies situate multi-agent LLMs within enterprise decision workflows that require legal accountability, cross-functional interpretation, and auditable reasoning (zhang2024proagent; 10.24963/ijcai.2024/890). Existing agent teams typically seek a single correct output, whereas organizational decisions often demand multiple perspectives, traceable rationales, and paths for mitigation when compliance risks are identified.
@@ 2.3. Human-AI Collaboration for Compliance Review
Regulatory compliance is a high-stakes, knowledge-intensive process traditionally handled by legal and domain experts. Recent work emphasizes that AI governance is sociotechnical: effective systems must integrate organizational workflows, stakeholder perspectives, and human expertise rather than focusing solely on technical accuracy (10.1145/3706599.3706747). Designing such systems also extends beyond interface design to model behavior and training data, requiring close collaboration between UX, engineering, and domain experts (subramonyom2022human).
AI tools have begun to assist in compliance-related work, but with mixed results. Studies show that GPT-4 can help developers prepare documentation for the EU AI Act (sovrano2025simplifying) by identifying gaps, though outputs were often incomplete or irrelevant and required expert intervention . Similarly, LLM-based assistants in legal practice can support case research or document review, yet even with retrieval augmentation, hallucination remains nontrivial (17–33%) (magesh2025hallucination). These results suggest AI can reduce workload, but expert oversight is indispensable.
@@ 2.4. Generative Agents for Critical Thinking and Advice
Generative agents extend LLM capabilities by storing experiences, reflecting on past interactions, and using these reflections to guide future planning. Such mechanisms improve task performance and produce more coherent long-term behavior (park2023generative; shinn2023reflexion). Beyond synthetic tasks, reflective techniques have also been adapted for coaching scenarios: a GPT-4-based chatbot using motivational interviewing improved users’ readiness for behavior change while mitigating potential harms through dialogue design (MEYER2025103514). These findings highlight the potential of agents to act as cognitive partners that encourage reflection rather than dispensing answers.
At the same time, reflective support raises new cognitive and interaction challenges. Studies emphasize that users collaborating with generative tools must develop self-awareness of their goals, calibrate confidence in AI capabilities, and flexibly adapt their strategies (10.1145/3613904.3642902). Experimental evidence further shows that the design of AI assistance profoundly shapes cognition and motivation: when AI generated ideas all at once, participants invested more effort in synthesizing suggestions, whereas incremental support led to higher confidence in their own ideas (10.1145/3706598.3713649). These results suggest that how AI support is structured directly influences user agency.
Building on this literature, recent work calls for positioning AI contributions to complement rather than dominate human reasoning (reichertsAIHelpMe2025). Yet prior studies remain focused on general creativity and coaching contexts, without addressing enterprise compliance workflows where reflective synthesis must reconcile legal, technical, and managerial perspectives. Our system fills this gap by embedding reflective generative agents into a roundtable framework, encouraging critical thinking and multi-perspective planning in regulated industry settings.
@@ 7.4. Data Collection
Quantitative Measures. We collected standardized usability and workload metrics through validated instruments. SUS (bangor2008empirical) provided 10-item Likert scale measurements (1-5) of system usability, with all questions rephrased as positive statements to ensure consistent interpretation. NASA-TLX (hart1988development) assessed cognitive workload across three key dimensions (Mental Demand, Effort, and Frustration) using the same 5-point scale format where higher scores indicate lower workload. Additionally, we measured trust and adoption metrics through custom 5-point Likert scales covering: Trust in Output, Usage Intention, Workflow Integration, and Explanation Clarity.
@@ 7.5. Data Analysis
Qualitative interview data were analyzed using thematic analysis (braun2006using), with two researchers independently coding transcripts to identify recurring themes related to system effectiveness, workflow integration, and areas for improvement.
@@ 9. Discussion
This study tackled the specific challenge of enterprise compliance review in the beauty tech industry, where AI-driven products face unique regulatory complexities around biometric data handling, skin-tone fairness, and cross-border compliance (Kelly_2025; 10.1145/3630106.3659038). We addressed three research questions: 1) current practices and challenges in beauty tech compliance review, 2) the design of a multi-agent LLM system to address these challenges, and 3) the effectiveness and user perception of such a system. We conducted a formative study with six experts (two legal, four IT) to identify domain-specific challenges. Based on these findings, we developed BeautyGuard, a multi-agent roundtable system that provides automated pre-review, role-specialized analysis, and graded risk mitigation. Six participants evaluated BeautyGuard using standardized metrics (SUS, NASA-TLX). The results showed that BeautyGuard achieved exceptional usability (SUS: 77.5/100), minimal workload across all dimensions.
While our focus was beauty tech compliance, the underlying challenges—fragmented pre-review processes, departmental silos, and binary decision-making—are common across regulated industries (10.1145/3706599.3706747; subramonyom2022human). This suggests that our multi-agent roundtable approach has broader applicability beyond the beauty domain, with potential applications in healthcare, financial services, and other regulated sectors that require coordinated expert review processes.
@@ 9.3. How to generalize the current approach from beauty tech to other regulated industries?
Our beauty tech focus revealed domain-specific challenges—biometric data handling, skin-tone fairness, and cultural sensitivity—that required specialized compliance knowledge (Kelly_2025; 10.1145/3630106.3659038). However, the underlying organizational challenges we identified are prevalent across regulated industries. Multi-agent systems have shown effectiveness in coordinating diverse expertise for complex decision-making tasks (hong2024metagpt; qian2023communicative), and our results demonstrate this applies to enterprise compliance contexts.
Similar principles can be generalized to other regulated domains since core challenges—fragmented processes, departmental silos, and binary decision-making—are documented across healthcare (subramonyom2022human), financial services, and other high-stakes industries (10.1145/3706599.3706747). The fundamental multi-agent architecture and roundtable coordination model remain applicable, with only domain-specific knowledge bases and regulatory mappings requiring adaptation.